# Activation of anomalous Hall effect and orbital magnetization by domain walls in altermagnets

### v1

@@ VII Acknowledgement
Appendix A Symmetry protection of Weyl nodal lines along $\Gamma$Z and MA
In this Appendix, we discuss the symmetry protection of the Weyl nodal lines along $\Gamma$Z and MA lines seen in the band structure in Fig.4(b) for the case of a uniform profile $\vec{N}=\hat{y}$. As will be shown below, the Weyl nodal line along MA is protected by a combination of a nonsymmorphic mirror $g_{1}=\{\mathcal{M}_{x}|\frac{1}{2}\frac{1}{2}\frac{1}{2}\}$ and an antiunitary symmetry $\tilde{g}_{2}=\Theta\{C_{2y}|\frac{1}{2}\frac{1}{2}\frac{1}{2}\}$ which involves the time reversal $\Theta$. (Tilde on $\tilde{g}_{2}$ denotes its anti-unitarity.) Meanwhile, the nodal line along $\Gamma$Z is not protected by any symmetry and can be gapped out by a symmetry-conforming pertubation.
Figure 9: Demonstration how the symmetry-conforming $\mathcal{H}_{\rm pert}$ gaps out the nodal line along $\Gamma$Z.
To see the absence of a symmetry protection for the nodal line along $\Gamma$Z, we supplement the following symmetry-conforming perturbation, which takes the form of a magnetic field along x-axis, into the model Hamiltonian $\mathcal{H}$ of Eq.(1)
@@ Appendix A Symmetry protection of Weyl nodal lines along $\Gamma$Z and MA
where $\eta$ is a coefficient. One can see that $\mathcal{H}_{\rm pert.}$ is symmetry-allowed from our discussion on the Landau theory for rutile altermagnets, where a Néel vector along y-axis induces a weak ferromagnetism with the magnetization along x-axis. Figure 9(b) shows how introducing $\mathcal{H}_{\rm pert}$ gaps out the nodal line along $\Gamma$Z in Fig.9(a). In contrast, the Weyl nodal line along MA is still protected.
@@
We will show that $\tilde{U}_{2}\psi_{\vec{k}}$ is also an eigenvector of $U_{1}$ but with the eigenvalue of $-e^{ik_{z}/2}$. To see that, we use Eq.(23):
@@ Appendix B Kubo formula for the domain wall Hall conductivity
where $\vec{q}=\vec{k}_{\parallel}=(k_{y},k_{z})^{T}$, $\gamma$ is a small broadening, |$\ket{u_{\vec{q}n}}$ denotes the Bloch eigenstates, and $v_{j}$ is the velocity operator defined below.
@@
[…]
$\approx$ $t_{ij}\psi^{\dagger}_{i}\psi_{j}\left[1+i\vec{r}_{ij}\cdot\frac{\vec{A}(\vec{r}_{i})+\vec{A}(\vec{r}_{j})}{2}\right],$
@@
$v_{l}(\vec{r}_{i})$ $=$ $\frac{\delta H[\vec{A}]}{\delta A_{l}(\vec{r}_{i})},$ (28)
@@ Appendix C Hall conductivity and orbital magnetization for collinear bulk
To compute the site-averaged Hall conductivity of Eq.(7) and a similar site-averaged orbital magnetization, we first compute the same quantities for the collinear bulk with the Néel vector orientation $\vec{N}_{i}$’s that appear in the DW Ansatz for $w_{\rm dw}=20$. Recall that $\vec{N}_{i}=\left(0,\sin\theta_{i},\cos\theta_{i}\right)^{T}$ and $\theta_{i}=\pi(2i+w_{\rm dw})/2w_{\rm dw}$ in the DW region $-w_{\rm dw}/2<i<w_{\rm dw}/2$. Figure 10 shows the $\mu$-dependence of the Hall conductivity $\sigma^{H}_{yz}(\vec{N}_{i})$ and that of the orbital magnetization $M_{\text{orb},x}(\vec{N}_{i})$. Note that $\vec{N}_{i=0}=\hat{y}$.

### v2

@@ VII Acknowledgement
[…]
@@
[…]
@@
We will show that $\tilde{U}_{2}\psi_{\vec{k}}$ is also an eigenvector of $U_{1}$ but with the eigenvalue of $e^{ik_{z}/2}$. To see that, we use Eq.(23):
@@ Appendix B Kubo formula for the domain wall Hall conductivity
[…]
@@
$t_{ij}\psi^{\dagger}_{i}\psi_{j}$ $\rightarrow$ $t_{ij}\psi^{\dagger}_{i}\psi_{j}\exp\left(-i\frac{e}{\hbar}\int_{\vec{r}_{i}}^{\vec{r}_{j}}d\vec{r}\cdot\vec{A}\right),$ (27)
$\approx$ $t_{ij}\psi^{\dagger}_{i}\psi_{j}\left[1-i\frac{e}{\hbar}\vec{r}_{ij}\cdot\frac{\vec{A}(\vec{r}_{i})+\vec{A}(\vec{r}_{j})}{2}\right],$
@@
$v_{l}(\vec{r}_{i})$ $=$ $-\frac{1}{e}\frac{\delta H[\vec{A}]}{\delta A_{l}(\vec{r}_{i})},$ (28)
@@
In the rest of the Appendix, we will apply our expressions to the case of the slab geometry and sketch the derivation of a useful form of $v_{l}$ Eq.(29) for the Kubo formula. We will first derive the tight-binding Hamiltonian in the Fourier space and then express $v_{l}$ in relation to the matrix elements of the Hamiltonian.
The tight-binding Hamiltonian on the slab geometry has the following form
$H$ $=$ $\sum_{ii^{\prime}ll^{\prime}\mu\mu^{\prime}}t_{il\mu,i^{\prime}l^{\prime}\mu^{\prime}}\psi^{\dagger}_{il\mu}\psi_{i^{\prime}l^{\prime}\mu^{\prime}}+\text{ hc.},$ (30)
where site indices $i$ and $i^{\prime}$ specify the in-plane position. The layer indices $l$ and $l^{\prime}$ specify the out-of-plane position. $\mu$ and $\mu^{\prime}$ are compact indices that specify the sublattices and the spin. We note that the position vector consists of out-of-plane and in-plane components: $\vec{r}=(r_{\perp},\vec{\mathfrak{r}})$. The in-plane translational invariance is encoded in the hopping integral $t_{il\mu,i^{\prime}l^{\prime}\mu^{\prime}}=t_{l\mu,l^{\prime}\mu^{\prime}}(\vec{\mathfrak{r}}_{i^{\prime}}-\vec{\mathfrak{r}}_{i})$, i.e. it depends only on the displacement vector $\vec{\mathfrak{r}}_{i^{\prime}}-\vec{\mathfrak{r}}_{i}$ within the slab plane. The Fourier form of $H$ is given by
$H$ $=$ $\sum_{\vec{q},ll^{\prime}\mu\mu^{\prime}}\psi^{\dagger}_{\vec{q}l\mu}\mathcal{H}_{l\mu,l^{\prime}\mu^{\prime}}(\vec{q})\psi_{\vec{q}l^{\prime}\mu^{\prime}},$ (31)
where
$\mathcal{H}_{l\mu,l^{\prime}\mu^{\prime}}(\vec{q})$ $=$ $t_{l\mu,l^{\prime}\mu^{\prime}}(\vec{q})+t_{l^{\prime}\mu^{\prime},l\mu}^{*}(\vec{q})$ (32)
$t_{l\mu,l^{\prime}\mu^{\prime}}(\vec{q})$ $=$ $\sum_{i^{\prime}}t_{l\mu,l^{\prime}\mu^{\prime}}(\vec{\mathfrak{r}}_{ii^{\prime}})e^{i\vec{q}\cdot\vec{\mathfrak{r}}_{ii^{\prime}}},$ (33)
where $\vec{\mathfrak{r}}_{ii^{\prime}}=\vec{\mathfrak{r}}_{i^{\prime}}-\vec{\mathfrak{r}}_{i}$. We have also used the Fourier transformation
$\psi_{il\mu}$ $=$ $\frac{1}{\sqrt{A}}\sum_{\vec{q}}\psi_{\vec{q}l\mu}e^{i\vec{q}\cdot\vec{\mathfrak{r}}_{i}}.$
AA is the area of the slab. The velocity operator Eq.(29) can be computed similarly. Following Eq.(29), we can write
$\vec{v}$ $=$ $\frac{1}{\hbar}\sum_{ii^{\prime}ll^{\prime}\mu\mu^{\prime}}it_{il\mu,i^{\prime}l^{\prime}\mu^{\prime}}(r_{\perp,ll^{\prime}}\hat{x}+\vec{\mathfrak{r}}_{ii^{\prime}})\psi_{il\mu}^{\dagger}\psi_{i^{\prime}l^{\prime}\mu^{\prime}}+\text{hc.}.$
Taking the same Fourier transformation akin to that for the Hamiltonian, we arrive at the following form
$v_{a}$ $=$ $\sum_{\vec{q},ll^{\prime}\mu\mu^{\prime}}\psi^{\dagger}_{\vec{q}l\mu}\mathcal{V}^{a}_{l\mu,l^{\prime}\mu^{\prime}}(\vec{q})\psi_{\vec{q}l^{\prime}\mu^{\prime}}.$ (35)
For the in-plane components $a=y,z,$
$\mathcal{V}^{a}_{l\mu,l^{\prime}\mu^{\prime}}(\vec{q})$ $=$ $\frac{1}{\hbar}\frac{\partial\mathcal{H}_{l\mu,l^{\prime}\mu}(\vec{q})}{\partial q_{a}},$ (36)
whereas the out-of-plane component $a=x$ corresponds to
$\mathcal{V}^{x}_{l\mu,l^{\prime}\mu^{\prime}}(\vec{q})$ $=$ $\frac{i}{\hbar}r_{\perp,ll^{\prime}}\mathcal{H}_{l\mu,l^{\prime}\mu^{\prime}}(\vec{q}).$ (37)
Note that there is no summation over the indices $l$ and $l^{\prime}$ in the right-hand side of the above equation. Finally, the notation for the matrix elements of the velocity operator in Eq.(29) is given by
[…]
@@ Appendix C Hall conductivity and orbital magnetization for collinear bulk
To compute the site-averaged Hall conductivity of Eq.(7) and a similar site-averaged orbital magnetization, we first compute the same quantities for the collinear bulk with the Néel vector orientation $\vec{N}_{i}$’s that appear in the DW Ansatz for $w_{\rm dw}=20$. Recall that $\vec{N}_{i}=\left(0,\sin\theta_{i},\cos\theta_{i}\right)^{T}$ and $\theta_{i}=\pi(2i+w_{\rm dw})/2w_{\rm dw}$ in the DW region $w_{\rm dw}/2<i<w_{\rm dw}/2$. Figure 10 shows the $\mu$-dependence of the Hall conductivity $\sigma^{H}_{yz}(\vec{N}_{i})$ and that of the orbital magnetization $M_{\text{orb},x}(\vec{N}_{i})$. Note that $\vec{N}_{i=0}=\hat{y}$.